OpenStudy (anonymous):

find the slope of the line tangent to the following curve where x = 1...

OpenStudy (anonymous):

\[y = \sin[x - \tan (\frac{ \Pi }{ 4 }x^{66})] + x ^{\frac{ 1 }{ 1 + 66\Pi }}\]

OpenStudy (anonymous):

if you cant see the exponent part at the end, its: \[\frac{ 1 }{ 1+66\Pi }\]

OpenStudy (anonymous):

answer is -99Pi

OpenStudy (anonymous):

im currently at this step: \[\frac{ -66\Pi }{ 4} + \frac{ 1 }{ 1 + 66\Pi }\]

zepdrix (zepdrix):

Oh goodness.. this one again iop? XD

OpenStudy (anonymous):

heh

zepdrix (zepdrix):

Hmm

zepdrix (zepdrix):

What is that symbol? Is that suppose to be pi?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Why don't u use lowercase? D: that's so confusing..

OpenStudy (anonymous):

i was wondering if that was confusing people

OpenStudy (anonymous):

i guess so

zepdrix (zepdrix):

Hmm you sure the step you got to is correct so far? :O

OpenStudy (anonymous):

i got this for y prime: \[y' = \cos[x - \tan(\frac{ \pi }{ 4 }x^{66})](1 - \sec^2(\frac{ \pi }{ 4 }x^{66}))(\frac{ 66\pi }{ 4 }x^{65}) + \frac{ 1 }{ 1+66\pi }x^\frac{ 1 }{ 1+66\pi }-1\]

zepdrix (zepdrix):

So what i tried to do is... After you get a common denominator and all that jazz.. you have something that looks like a polynomial in the numerator. And I'm not seeing it factor nicely :( at least not in any way that will cancel with the denominator. Hmmm

OpenStudy (anonymous):

plugged 1 in.. \[y'(1) = 1(-1)(\frac{ 66\pi }{ 4 })(\frac{ 1 }{ 1+ 66\pi })\]

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

the factoring is the confusing part... how am i supposed to get it to -99pi

OpenStudy (anonymous):

unless the steps i did to get it to that point were incorrect...i dont think so

zepdrix (zepdrix):

 Hmm yah your steps look good so far :o